# How Private is DP-SGD?

### v1

@@ 2.2 Adjacency Notions
Datasets $\bm{x},\bm{x}^{\prime}\in\mathcal{X}^{*}$ are said to be add-remove adjacent if there exists an $i$ such that $\bm{x}^{\prime}=\bm{x}_{-i}$ or vice-versa (where $\bm{x}_{-i}$ represents the dataset obtained by removing the $i$-th record in $\bm{x}$).
@@
The privacy analysis of DP-SGD is typically done for the Poisson batch sampler $\mathcal{P}$ (Abadi et al., 2016; Mironov, 2017), with respect to the add-remove adjacency. However, it is impossible to analyze the privacy of $\textsc{ABLQ}_{\mathcal{D}}$ or $\textsc{ABLQ}_{\mathcal{S}}$ with respect to the add-remove adjacency because the batch samplers $\mathcal{D}$ and $\mathcal{S}$ require that the number of records $n$ equals $b\cdot T$. On the other hand, using the substitution adjacency for $\mathcal{D}$ and $\mathcal{S}$ leads to an unfair comparison to $\textsc{ABLQ}_{\mathcal{P}}$ whose analysis is with respect to the add-remove adjacency. Thus, to make a fair comparison, we consider the following adjacency (considered by Kairouz et al. (2021)).
@@ 2.3 Hockey Stick Divergence
For all $\varepsilon\in\mathbb{R}$, the $e^{\varepsilon}$-hockey stick divergence between $P$ and $Q$ is $D_{e^{\varepsilon}}(P\|Q):=\sup_{E}P(E)-e^{\varepsilon}Q(E)$.
It is immediate to see that $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-$\mathsf{DP}$ iff for all adjacent $\bm{x}\sim\bm{x}^{\prime}$, it holds that $D_{e^{\varepsilon}}(\mathcal{M}(\bm{x})||\mathcal{M}(\bm{x}^{\prime}))\leq\delta$.
@@
If $(P,Q)$ dominates $\mathcal{M}$, then for all $\varepsilon\geq 0$, $\delta_{\mathcal{M}}(\varepsilon)\leq\max\{D_{e^{\varepsilon}}(P\|Q),D_{e^{\varepsilon}}(Q\|P)\}$. We say that $(P,Q)$ tightly dominates a mechanism $\mathcal{M}$ if $(P,Q)$ dominates $\mathcal{M}$ and there exist adjacent datasets $\bm{x}\to_{z}\bm{x}^{\prime}$ such that $D_{e^{\varepsilon}}(P\|Q)=D_{e^{\varepsilon}}(\mathcal{M}(\bm{x})\|\mathcal{M}(\bm{x}^{\prime}))$ holds for all $\varepsilon\in R$ (note that this includes $\varepsilon<0$); in this case, $\delta_{\mathcal{M}}(\varepsilon)=\max\{D_{e^{\varepsilon}}(P\|Q),D_{e^{\varepsilon}}(Q\|P)\}$. Thus, tightly dominating pairs completely characterize the privacy loss of a mechanism (although they are not guaranteed to exist for all mechanisms.) Dominating pairs behave nicely under mechanism compositions. Namely, if $(P_{1},Q_{1})$ dominates $\mathcal{M}_{1}$ and $(P_{2},Q_{2})$ dominates $\mathcal{M}_{2}$, then $(P_{1}\otimes P_{2},Q_{1}\otimes Q_{2})$ dominates the (adaptively) composed mechanism $\mathcal{M}_{1}\circ\mathcal{M}_{2}$.
@@ 3 Dominating Pairs for $\textsc{ABLQ}_{\mathcal{B}}$
Zhu et al. (2022) showed that the tightly dominating pair for a single step of $\textsc{ABLQ}_{\mathcal{P}}$, a Poisson sub-sampled Gaussian mechanism, is given by the pair $(A=(1-q)\mathcal{N}(0,\sigma^{2})+q\mathcal{N}(1,\sigma^{2}),B=\mathcal{N}(0,\sigma^{2}))$, where $q$ is the sub-sampling probability of each record, namely $q=b/n$, and in the case where $n=b\cdot T$, we have $q=1/T$. Since $\textsc{ABLQ}_{\mathcal{P}}$ is a $T$-fold composition of this Poisson subsampled Gaussian mechanism, it follows that the tightly dominating pair for $\textsc{ABLQ}_{\mathcal{P}}$ is $(P_{\mathcal{P}}:=A^{\otimes T},Q_{\mathcal{P}}:=B^{\otimes T})$.
The hockey stick divergence $D_{e^{\varepsilon}}(P_{\mathcal{P}}\|Q_{\mathcal{P}})$ does not have a closed-form expression, but there are privacy accountants based on the methods of Rényi differential privacy (RDP)  (Mironov, 2017) as well as privacy loss distributions (PLD)  (Meiser & Mohammadi, 2018; Sommer et al., 2019), the latter providing numerically accurate algorithms (Koskela et al., 2020; Gopi et al., 2021; Ghazi et al., 2022; Doroshenko et al., 2022), and have been the basis of multiple open-source implementations from both industry and academia including Prediger & Koskela (2020); Google’s DP Library. (2020); Microsoft. (2021). While Rényi-DP based accounting provides an upper bound on $\max\{D_{e^{\varepsilon}}(P_{\mathcal{P}}\|Q_{\mathcal{P}}),D_{e^{\varepsilon}}(Q_{\mathcal{P}}\|P_{\mathcal{P}})\}$, the PLD based accounting implementations can provide upper and lower bounds on $\max\{D_{e^{\varepsilon}}(P_{\mathcal{P}}\|Q_{\mathcal{P}}),D_{e^{\varepsilon}}(Q_{\mathcal{P}}\|P_{\mathcal{P}})\}$ to high accuracy, as controlled by a certain discretization parameter.
@@
In this case, it is easy to see that the distributions $A=\textsc{ABLQ}_{\mathcal{S}}(\bm{x})$ and $B=\textsc{ABLQ}_{\mathcal{S}}(\bm{x}^{\prime})$ are given as:
@@
where $\mathbf{1}$ denotes the all-$1$’s vector in $\mathbb{R}^{T}$ and $e_{t}$ denotes the $t$-th standard basis vector in $\mathbb{R}^{T}$. Shifting the distributions by $b\cdot\mathbf{1}$ does not change the hockey stick divergence $D_{e^{\varepsilon}}(A\|B)$, hence we might as well consider the pair
$P_{\mathcal{S}}$ $:=~{}\textstyle\sum_{t=1}^{T}\frac{1}{T}\cdot\mathcal{N}(2e_{t},\sigma^{2}I)$ (1)
$Q_{\mathcal{S}}$ $:=~{}\textstyle\sum_{t=1}^{T}\frac{1}{T}\cdot\mathcal{N}(e_{t},\sigma^{2}I)$ (2)
@@ 4.1 $\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{S}}$
For all $\varepsilon\geq 0$, it holds that $\delta_{\mathcal{S}}(\varepsilon)\leq\delta_{\mathcal{D}}(\varepsilon)$.
@@ 4.2 $\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{P}}$
There exist $\varepsilon_{0},\varepsilon_{1}\geq 0$ such that,
@@
    $\forall\varepsilon\in[0,\varepsilon_{0})$: it holds that $\delta_{\mathcal{D}}(\varepsilon)>\delta_{\mathcal{P}}(\varepsilon)$, and
@@
    $\forall\varepsilon>\varepsilon_{1}$: it holds that $\delta_{\mathcal{D}}(\varepsilon)<\delta_{\mathcal{P}}(\varepsilon)$.
We use the Gaussian measure of a halfspace.
Proposition 4.3.

### v2

@@ 2.2 Adjacency Notions
Datasets $\bm{x},\bm{x}^{\prime}\in\mathcal{X}^{*}$ are said to be add-remove adjacent if there exists an $i$ such that $\bm{x}^{\prime}=\bm{x}_{-i}$ or vice-versa (where $\bm{x}_{-i}$ represents the dataset obtained by removing the $i$th record in $\bm{x}$).
@@
The privacy analysis of $\mathsf{DP}\text{-}\mathsf{SGD}$ is typically done for the Poisson batch sampler $\mathcal{P}$ (Abadi et al., 2016; Mironov, 2017), with respect to the add-remove adjacency. However, it is impossible to analyze the privacy of $\textsc{ABLQ}_{\mathcal{D}}$ or $\textsc{ABLQ}_{\mathcal{S}}$ with respect to the add-remove adjacency because the batch samplers $\mathcal{D}$ and $\mathcal{S}$ require that the number of records $n$ equals $b\cdot T$. On the other hand, using the substitution adjacency for $\mathcal{D}$ and $\mathcal{S}$ leads to an unfair comparison to $\textsc{ABLQ}_{\mathcal{P}}$ whose analysis is with respect to the add-remove adjacency. Thus, to make a fair comparison, we consider the following adjacency (proposed by Kairouz et al. (2021)).
@@
We note that our separations where $\varepsilon_{\mathcal{P}}(\delta)\ll\varepsilon_{\mathcal{S}}(\varepsilon)$, such as in Figure 1, also hold under the substitution adjacency, by using “group privacy”, namely if a mechanism satisfies $(\varepsilon,\delta)$-$\mathsf{DP}$ with respect to zero-out adjacency, then it satisfies $(2\varepsilon,\delta\cdot(1+e^{\varepsilon}))$-$\mathsf{DP}$ (see, e.g., (Vadhan, 2017)).
@@ 2.3 Hockey Stick Divergence
For all $\varepsilon\in\mathbb{R}$, the $e^{\varepsilon}$-hockey stick divergence between $P$ and $Q$ is $D_{e^{\varepsilon}}(P\|Q):=\sup_{E}\{P(E)-e^{\varepsilon}Q(E)\}$.
It is immediate to see that $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-$\mathsf{DP}$ iff for all adjacent $\bm{x}\sim\bm{x}^{\prime}$, it holds that $D_{e^{\varepsilon}}(\mathcal{M}(\bm{x})\|\mathcal{M}(\bm{x}^{\prime}))\leq\delta$.
@@
If $(P,Q)$ dominates $\mathcal{M}$, then for all $\varepsilon\geq 0$, $\delta_{\mathcal{M}}(\varepsilon)\leq\max\{D_{e^{\varepsilon}}(P\|Q),D_{e^{\varepsilon}}(Q\|P)\}$. We say that $(P,Q)$ tightly dominates a mechanism $\mathcal{M}$ if $(P,Q)$ dominates $\mathcal{M}$ and there exist adjacent datasets $\bm{x}\to_{z}\bm{x}^{\prime}$ such that $D_{e^{\varepsilon}}(P\|Q)=D_{e^{\varepsilon}}(\mathcal{M}(\bm{x})\|\mathcal{M}(\bm{x}^{\prime}))$ holds for all $\varepsilon\in R$ (note that this includes $\varepsilon<0$); in this case, $\delta_{\mathcal{M}}(\varepsilon)=\max\{D_{e^{\varepsilon}}(P\|Q),D_{e^{\varepsilon}}(Q\|P)\}$. Thus, tightly dominating pairs completely characterize the privacy loss of a mechanism (although they are not guaranteed to exist for all mechanisms). [^5] Zhu et al. (2022) define “tightly dominating pair” differently, in a manner that is guaranteed to exist. They additional define the notion of a “worst-case pair”, which is a pair of adjacent datasets $\bm{x}\sim\bm{x}^{\prime}$ such that $(\mathcal{M}(\bm{x}),\mathcal{M}(\bm{x}^{\prime}))$ is a tightly dominating pair. Thus, our notion of “tightly dominating pair” refers precisely to the pair $(\mathcal{M}(\bm{x}),\mathcal{M}(\bm{x}^{\prime}))$ for a worst-case adjacent pair $\bm{x}$, $\bm{x}^{\prime}$. It is also worth noting that our notation for “tightly dominating pairs” is asymmetric as it only considers pairs $\bm{x}\to_{z}\bm{x}^{\prime}$; the reverse setting is handled implicitly because $(P,Q)$ dominates $(\mathcal{M}(\bm{x}),\mathcal{M}(\bm{x}^{\prime}))$ if and only if $(Q,P)$ dominates $(\mathcal{M}(\bm{x}^{\prime}),\mathcal{M}(\bm{x}))$. Dominating pairs behave nicely under mechanism compositions. Namely, if $(P_{1},Q_{1})$ dominates $\mathcal{M}_{1}$ and $(P_{2},Q_{2})$ dominates $\mathcal{M}_{2}$, then $(P_{1}\otimes P_{2},Q_{1}\otimes Q_{2})$ dominates the (adaptively) composed mechanism $\mathcal{M}_{1}\circ\mathcal{M}_{2}$.
@@ 3 Dominating Pairs for $\textsc{ABLQ}_{\mathcal{B}}$
Zhu et al. (2022) showed [^6] This was implicit in prior work, e.g., (Koskela et al., 2020). that the tightly dominating pair for a single step of $\textsc{ABLQ}_{\mathcal{P}}$, a Poisson sub-sampled Gaussian mechanism, is given by the pair $(A=(1-q)\mathcal{N}(0,\sigma^{2})+q\mathcal{N}(1,\sigma^{2}),B=\mathcal{N}(0,\sigma^{2}))$, where $q$ is the sub-sampling probability of each record, namely $q=b/n$, and in the case where $n=b\cdot T$, we have $q=1/T$. Since $\textsc{ABLQ}_{\mathcal{P}}$ is a $T$-fold composition of this Poisson subsampled Gaussian mechanism, it follows that the tightly dominating pair for $\textsc{ABLQ}_{\mathcal{P}}$ is $(P_{\mathcal{P}}:=A^{\otimes T},Q_{\mathcal{P}}:=B^{\otimes T})$.
The hockey stick divergence $D_{e^{\varepsilon}}(P_{\mathcal{P}}\|Q_{\mathcal{P}})$ does not have a closed-form expression, but there are privacy accountants based on the methods of Rényi DP (RDP)  (Mironov, 2017) as well as privacy loss distributions (PLD)  (Meiser & Mohammadi, 2018; Sommer et al., 2019), the latter providing numerically accurate algorithms (Koskela et al., 2020; Gopi et al., 2021; Ghazi et al., 2022; Doroshenko et al., 2022), and have been the basis of multiple open-source implementations from both industry and academia including (Prediger & Koskela, 2020; Google’s DP Library., 2020; Microsoft., 2021). While Rényi-DP-based accounting provides an upper bound on $\max\{D_{e^{\varepsilon}}(P_{\mathcal{P}}\|Q_{\mathcal{P}}),D_{e^{\varepsilon}}(Q_{\mathcal{P}}\|P_{\mathcal{P}})\}$, the PLD-based accounting implementations can provide upper and lower bounds on $\max\{D_{e^{\varepsilon}}(P_{\mathcal{P}}\|Q_{\mathcal{P}}),D_{e^{\varepsilon}}(Q_{\mathcal{P}}\|P_{\mathcal{P}})\}$ to high accuracy, as controlled by a certain discretization parameter.
@@
In this case, when the differing record falls in batch $t$, then output of the mechanism on all batches $t^{\prime}\neq t$ is centered at $-b$, and is centered at $-b+2$ (on input $\bm{x}$) or at $-b+1$ (on input $\bm{x}^{\prime}$) on batch $t$. Thus it follows that the distributions $A=\textsc{ABLQ}_{\mathcal{S}}(\bm{x})$ and $B=\textsc{ABLQ}_{\mathcal{S}}(\bm{x}^{\prime})$ are given as:
@@
where $\mathbf{1}$ denotes the all-$1$’s vector in $\mathbb{R}^{T}$ and $e_{t}$ denotes the $t$th standard basis vector in $\mathbb{R}^{T}$. Shifting the distributions by $b\cdot\mathbf{1}$ does not change the hockey stick divergence $D_{e^{\varepsilon}}(A\|B)$, hence we might as well consider the pair
$P_{\mathcal{S}}$ $:=~{}\textstyle\sum_{t=1}^{T}\frac{1}{T}\cdot\mathcal{N}(2e_{t},\sigma^{2}I),$ (1)
$Q_{\mathcal{S}}$ $:=~{}\textstyle\sum_{t=1}^{T}\frac{1}{T}\cdot\mathcal{N}(e_{t},\sigma^{2}I).$ (2)
@@ 4.1 $\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{S}}$
For all $\sigma,\varepsilon\geq 0$ and $T\geq 1$: $\delta_{\mathcal{S}}(\varepsilon)\leq\delta_{\mathcal{D}}(\varepsilon)$.
@@ 4.2 $\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{P}}$
For all $\sigma>0$ and $T>1$, there exist $\varepsilon_{0},\varepsilon_{1}\geq 0$ such that,
@@
    $\forall\varepsilon\in[0,\varepsilon_{0})$, it holds that $\delta_{\mathcal{D}}(\varepsilon)>\delta_{\mathcal{P}}(\varepsilon)$, and
@@
    $\forall\varepsilon>\varepsilon_{1}$, it holds that $\delta_{\mathcal{D}}(\varepsilon)<\delta_{\mathcal{P}}(\varepsilon)$.
We defer the detailed proof to Appendix C, and provide a proof sketch here. Part (a) is shown by first establishing that the total variation distance (corresponds to $D_{1}(\cdot\|\cdot)$) between $P_{\mathcal{D}}$ and $Q_{\mathcal{D}}$ is strictly larger than the total variation distance between $P_{\mathcal{P}}$ and $Q_{\mathcal{P}}$ when $T>1$ and $\sigma>0$. This implies that, $\delta_{\mathcal{D}}(0)>\delta_{\mathcal{P}}(0)$. By using the continuity of $D_{e^{\varepsilon}}(\cdot\|\cdot)$ in $\varepsilon$, we conclude the same for all $\varepsilon<\varepsilon_{0}$.
For part (b), we construct an explicit set $E$ such that $P_{\mathcal{P}}(E)-e^{\varepsilon}Q_{\mathcal{P}}(E)>\delta_{\mathcal{D}}(\varepsilon)$. In particular, we choose a halfspace $E:=\left\{w\in\mathbb{R}^{T}\big{|}\sum_{i}w_{i}>(\varepsilon+\log 2+T\log T)\sigma^{2}+{\frac{T}{2}}\right\}$ and show that $P_{\mathcal{P}}(E)-e^{\varepsilon}Q_{\mathcal{P}}(E)$ is at least $\frac{1}{2}\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}}-\frac{(T\log T+\log 2)\sigma}{\sqrt{T}}-\frac{\sqrt{T}}{2\sigma}\right)$. For large $\varepsilon$, the dominant term is $-\varepsilon\sigma/\sqrt{T}$. On other hand, $\delta_{\mathcal{D}}(\varepsilon)$ is at most $\Phi(-\varepsilon\sigma+\frac{1}{2\sigma})$ (from Proposition 3.1), which has the dominant term $-\varepsilon\sigma$. Since $-\varepsilon\sigma/\sqrt{T}$ decays slower than $-\varepsilon\sigma$, we get that for sufficiently large $\varepsilon_{1}$, it holds that $\delta_{\mathcal{D}}(\varepsilon)<\delta_{\mathcal{P}}(\varepsilon)$ for all $\varepsilon>\varepsilon_{1}$.
Conjecture 4.3.
Theorem 4.2 holds for $\varepsilon_{0}=\varepsilon_{1}$.
We do not rely on this conjecture being true in the rest of this paper. We provide a numerical example that validates Theorem 4.2 and provides evidence for 4.3. In Figure 2, for $\sigma=0.3$ and $T=10$, we plot the numerically computed $\delta_{\mathcal{D}}(\varepsilon)$ (using Proposition 3.1), as well as lower and upper bounds on $\delta_{\mathcal{P}}(\varepsilon)$, computed using the open source dp_accounting library (Google’s DP Library., 2020).
4.3 $\textsc{ABLQ}_{\mathcal{P}}$ vs $\textsc{ABLQ}_{\mathcal{S}}$
$D(E_{C})$ $\textstyle~{}=~{}1-D(\mathbb{R}^{T}\smallsetminus E_{C})$
$\textstyle~{}=~{}1-\prod_{t=1}^{T}\Phi\left(\frac{C-\mu_{t}}{\sigma}\right).$
$\textstyle~{}=~{}1-\Phi\left(\frac{C-2}{\sigma}\right)\cdot\Phi\left(\frac{C}{\sigma}\right)^{T-1}.$
Thus, we use the following lower bound:
In Figure 3, we set $\sigma=0.4$ and number of steps $T=10,000$ and plot $\delta_{\mathcal{D}}(\varepsilon)$, an upper bound on $\delta_{\mathcal{P}}(\varepsilon)$ (obtained using dp_accounting) and a lower bound on $\delta_{\mathcal{S}}(\varepsilon)$ as obtained via (3). We find that while $\delta_{\mathcal{P}}(4)\leq 1.18\cdot 10^{-5}$, $\delta_{\mathcal{S}}(4)\geq 0.226$, that is close to $\delta_{\mathcal{D}}(4)\approx 0.244$. Even $\delta_{\mathcal{S}}(12)\geq 7.5\cdot 10^{-5}$ is larger than $\delta_{\mathcal{P}}(4)$. While the $(4,1.2\cdot 10^{-5})$-$\mathsf{DP}$ guarantee of $\textsc{ABLQ}_{\mathcal{P}}$ could have been considered as sufficiently private, $\textsc{ABLQ}_{\mathcal{S}}$ only satisfies much worse privacy guarantees. We provide additional examples in Appendix A.
4.3.1 Intuition for 3.2
  • $\bm{x}=(x_{1}=-L,\ldots,x_{n-1}=-L,x_{n}=1)$, and
  • $\bm{x}^{\prime}=(x_{1}=-L,\ldots,x_{n-1}=-L,x_{n}=\bot)$.
The case of $L>1$ is not valid, since in this case $|\psi_{t}(x)|=L>1$. However, we can still ask how well the privacy of the $n$th example is preserved by $\textsc{ABLQ}_{\mathcal{B}}$, by considering the hockey stick divergence between $\textsc{ABLQ}_{\mathcal{B}}(\bm{x})$ and $\textsc{ABLQ}_{\mathcal{B}}(\bm{x}^{\prime})$.
In summary, the main differing aspect about $\textsc{ABLQ}_{\mathcal{S}}$ and $\textsc{ABLQ}_{\mathcal{P}}$ is that in the former, the non-differing examples can leak information about the location of the differing example in the shuffled order, but it is not the case in the latter. While we sketched an argument that works asymptotically as $L\to\infty$, we see glimpses of it already at $L=1$.
Our intuitive reasoning behind 3.2 is that even in the case of (vector-valued) adaptive query methods, in order to “leak the most information” about the differing record $x_{i}$ between $\bm{x}$ and $\bm{x}^{\prime}$, it seems natural that the query $\psi_{t}(x_{j})$ should evaluate to $-\psi_{t}(x_{i})$ for all $j\neq i$. If the query method satisfies this condition for all $t$, then it is easy to show that $(P_{\mathcal{S}},Q_{\mathcal{S}})$ tightly dominates $(\textsc{ABLQ}_{\mathcal{S}}(\bm{x}),\textsc{ABLQ}_{\mathcal{S}}(\bm{x}^{\prime}))$. 3.2 then asserts that this is indeed the worst case.
5 Conclusion & Future Directions
Several interesting directions remain to be investigated. In our work, we provide a technique to only provide a lower bound on the privacy guarantee when using a shuffle batch sampler. It would be interesting to have a tight accounting method for $\textsc{ABLQ}_{\mathcal{S}}$. A first step towards this could be to establish 3.2, which if true, might make numerical accountants for computing the hockey stick divergence possible. While this involves computing a high-dimensional integral, it might be possible to approximate using importance sampling; e.g., such approaches have been attempted for $\textsc{ABLQ}_{\mathcal{P}}$ (Wang et al., 2023). Also, our approach for analyzing the privacy with shuffle batch sampler is limited to a “single epoch” mechanism, whereas in practice, it is common to use $\mathsf{DP}\text{-}\mathsf{SGD}$ with multiple epochs. Extending our approach to multiple epochs will be interesting.
However, it remains to be seen how the utility of $\mathsf{DP}\text{-}\mathsf{SGD}$ would be affected when we use the correct privacy analysis for $\textsc{ABLQ}_{\mathcal{S}}$ instead of the analysis for $\textsc{ABLQ}_{\mathcal{P}}$, which has been used extensively so far and treated as a good “approximation”. Alternative approaches such as $\mathsf{DP}\text{-}\mathsf{FTRL}$ (Kairouz et al., 2021; McMahan et al., 2022) that do not rely on amplification might turn out to be better if we instead use the correct analysis for $\textsc{ABLQ}_{\mathcal{S}}$, in the regimes where such methods are currently reported to under perform compared to $\mathsf{DP}\text{-}\mathsf{SGD}$.
An important point to note is that the model of shuffle batch sampler we consider here is a simple one. There are various types of data loaders used in practice, which are not necessarily captured by our model. For example tf.data.Dataset.shuffle takes in a parameter of buffer size $b$. It returns a random record among the first $b$ records, and immediately replaces it with the next record ($(b+1)$st in this case), and continues repeating this process. This leads to an asymmetric form of shuffling, when the dataset size exceeds the size of the buffer. Such batch samplers might thus require more careful privacy analysis.
The notion of $\mathsf{DP}$ aims to guarantee privacy even in the worst case. For example in the context of $\mathsf{DP}\text{-}\mathsf{SGD}$, it aims to protect privacy of a single record even when the model trainer and all other records participating in the training are colluding to leak information about this one record. And moreover, releasing the final model is assumed to leak as much information as releasing all intermediate iterates. Such strong adversarial setting might make obtaining good utility to be difficult. Alternative techniques for privacy amplification such as amplification by iteration (Feldman et al., 2018; Altschuler & Talwar, 2022) or through convergence of Langevin dynamics (Chourasia et al., 2021) have been studied, where only the last iterate of the model is released. However, these analyses rely on additional assumptions such as convexity and smoothness of the loss functions. Investigating whether it is possible to relax these assumptions to make amplification by iteration applicable to non-convex models, even if under some assumptions that are applicable to the ones used in practice, is an interesting future direction.
Acknowledgments
We thank the anonymous ICML reviewers for their thoughtful comments and suggestions which have improved the presentation of this paper. We also thank Christian Janos Lebeda, Matthew Regehr, Gautam Kamath, and Thomas Steinke for correspondence about their concurrent work (Lebeda et al., 2024).
Impact Statement
This paper presents work that points out that the particular type of batch sampling used can play a significant role in the privacy analysis of DP-SGD type of algorithms. The impact we see coming from our work is to make practitioners of DP more aware of these gaps. There are many potential societal consequences of DP itself, none which we feel must be specifically highlighted here.
Appendix A Additional Evaluations
A.1 $\varepsilon$ vs. $\sigma$ for fixed $\delta$ and $T$
We first plot $\varepsilon$ against $\sigma$ for fixed $\delta$ and $T$. We compute upper bounds on $\varepsilon_{\mathcal{P}}(\delta)$ using Rényi DP (RDP) as well as using privacy loss distributions (PLD). These accounting methods are provided in the open source Google dp_accounting library (Google’s DP Library., 2020).
  • In Figure 4, we set $\delta=10^{-6}$ and number of steps $T=100,000$. In particular, for $\sigma=0.4$, we find that $\varepsilon_{\mathcal{P}}(\delta)\leq 3$ (as per PLD accounting) and $\varepsilon_{\mathcal{P}}(\delta)\leq 4.71$ (as per RDP accounting), whereas on the other hand, $\varepsilon_{\mathcal{S}}(\delta)\geq 14.45$, which is very close to $\varepsilon_{\mathcal{D}}(\delta)$. For $\sigma=1.3$, we find that $\varepsilon_{\mathcal{P}}(\delta)<0.01$ (as per PLD accounting), whereas, $\varepsilon_{\mathcal{S}}(\delta)>0.029$.
  • In Figure 5, we set $\delta=10^{-5}$ and number of steps $T=1000$. In particular, for $\sigma=0.7$, $\varepsilon_{\mathcal{P}}(\delta)\leq 0.61$ (as per PLD accounting) and $\varepsilon_{\mathcal{P}}(\delta)\leq 1.64$ (as per RDP accounting), whereas on the other hand, $\varepsilon_{\mathcal{S}}(\delta)\geq 6.528$ and $\varepsilon_{\mathcal{D}}(\delta)\approx 6.652$. For $\sigma=1.3$, we find that $\varepsilon_{\mathcal{P}}(\delta)<0.092$ (as per PLD accounting), whereas, $\varepsilon_{\mathcal{S}}(\delta)>0.83$.
A.2 $\delta$ vs. $\varepsilon$ for fixed $\sigma$ and $T$
Next, we plot $\delta$ against $\varepsilon$ for fixed $\sigma$ and $T$. We compute upper bounds on $\delta_{\mathcal{P}}(\varepsilon)$ using Rényi DP (RDP) as well as using privacy loss distributions (PLD).
Appendix B Proof of Theorem 4.1: ($\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{S}}$)
We use the joint convexity property of hockey stick divergence. While this is standard, we include a proof for completeness.
Lemma B.1 (Joint Convexity of Hockey Stick Divergence).
Given two mixture distributions $P=\sum_{i=1}^{m}\alpha_{i}P_{i}$ and $Q=\sum_{i=1}^{m}\alpha_{i}Q_{i}$, it holds for all $\varepsilon\in\mathbb{R}$ that $D_{e^{\varepsilon}}(P\|Q)~{}\leq~{}\sum_{i}\alpha_{i}D_{e^{\varepsilon}}(P_{i}\|Q_{i})\,.$
Proof.
We have that
$D_{e^{\varepsilon}}(P\|Q)$ $~{}=~{}\sup_{E}\left\{P(E)-e^{\varepsilon}Q(E)\right\}$
$~{}=~{}\sup_{E}\left\{\sum_{i=1}^{m}\alpha_{i}\left(P_{i}(E)-e^{\varepsilon}Q_{i}(E)\right)\right\}$
$~{}\leq~{}\sum_{i=1}^{m}\alpha_{i}\cdot\sup_{E}\left\{P_{i}(E)-e^{\varepsilon}Q_{i}(E)\right\}~{}=~{}\sum_{i=1}^{m}\alpha_{i}D_{e^{\varepsilon}}(P_{i}\|Q_{i})\,.\qed$
Thus, it follows that shuffling the dataset first cannot degrade the privacy guarantee of any mechanism as shown below.
Lemma B.2.
Proof.
Since $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-$\mathsf{DP}$ it follows that $D_{e^{\varepsilon}}(P_{\pi}\|Q_{\pi})\leq\delta$ for all permutations $\pi$. Thus, from Lemma B.1, it follows that $D_{e^{\varepsilon}}(P\|Q)\leq\sum_{\pi}\frac{1}{n!}D_{e^{\varepsilon}}(P_{\pi}\|Q_{\pi})\leq\delta$. Hence $\mathcal{M}_{\mathcal{S}}$ also satisfies $(\varepsilon,\delta)$-$\mathsf{DP}$. ∎
Proof of Theorem 4.1.
Appendix C Proof of Theorem 4.2 ($\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{P}}$)
We first state and prove some intermediate statements required for the proof of Theorem 4.2. We use the Gaussian measure of a halfspace.
Proposition C.1.